# LORE: A Large Generative Model for Search Relevance

### v1

@@ Abstract
Achievement. We introduce LORE(Large Generative Model for Relevance), a complete and sustainable framework of iterative practices for large language models(LLMs) in e-commerce search relevance, which achieves a cumulative +27% improvement on the online GoodRate metric. Over the past three years, this project has demonstrated significant improvements in relevance judgment and has undergone three full-scale iterations across key dimensions, including data, features, training paradigms, evaluation, and application. Throughout the iterative development of LORE, we have gained valuable experience and insights that we believe are worth sharing with the community in this report.
@@
Second, guided by our structural analysis, we propose a sophisticated two-stage training paradigm that uses progressive CoT synthesis to inject comprehensive capabilities via SFT, followed by a carefully designed RL phase for human preference alignment, and share insights from our comprehensive exploration of the associated training strategies.
@@
Finally, to overcome the challenges of real-time computation, we designed a query frequency-stratified deployment strategy that comprehensively transfers the offline LLM’s ability to online system, leading to substantial online performance gains. The complete methodology and findings provide valuable insights for post-training work in other vertical domains.
LORE serves as both a practical solution for developing and operationalizing advanced relevance systems in e-commerce and a methodological reference for post-training work across other vertical domains.
@@ 1 Introduction
Recent advances(walmartllm; ELLM; LREF; JD2; TaoSR1) have progressively transitioned the post-training paradigm from rudimentary classification-oriented Supervised Fine-Tuning (SFT) to sophisticated enhancement of reasoning capabilities. Although existing work has modeled chain-of-thought(CoT)(CoT) processes from multiple perspectives, notable gaps remain that limit comprehensive treatment of complex relevance tasks. ELLM(ELLM) conceptualizes the relevance task as a process of attribute extraction and matching between queries and items. However, this approach exhibits limited capability when handling boundary cases that require rule-based judgment. For instance, when matching the query ”lake blue top” with the item ”sky blue top,” the method lacks color discrimination rules and struggles to make accurate judgments. LREF(LREF) and TaoSR1(TaoSR1) not only focus on attribute matching but also explicitly model rule-following chains of thought, endowing the model with rule-awareness capabilities. Nevertheless, their failure to incorporate visual information results in blind spots in discrimination. Taking the query ”blue top” as an example, when the item title lacks color description but the item image clearly displays a blue top, the model fails to make correct judgments due to the absence of image information. These limitations stem from neglecting a fundamental question: What constitutes a comprehensive CoT representation for relevance tasks? This question probes the capability requirements for relevance assessment.
@@
To systematically address this question, we conducted an in-depth deconstructive analysis of the relevance task, detailed in Section 2. We decompose the relevance discrimination process into two stages: (1) Path Construction: mapping query requirements to specific attributes and establish semantic paths between queries and products; (2) Path Following: execute concrete discrimination at the attribute-product level based on the constructed paths. Furthermore, the model capabilities required throughout this process can be summarized as follows: (1) Knowledge and Reasoning Ability(deepseek; skywork; kimi1.5; knowledge1; knowledge2). The model relies on knowledge to parse domain-specific expressions in queries and obscure item information. Reasoning capability then builds upon this foundation, enabling the model to infer the true intent of queries and the derived attributes of items. As illustrated in Fig. 1(a), the model must first rely on knowledge to determine that ”morning C, evening A” refers to a combination of Vitamin C and retinol, and then deduce through reasoning that ”for mom” implies suitability for middle-aged women. (2) Multi-modal Matching Ability(multi-modal1; multi-modal2; multi-modal3).Item attributes are presented across multiple modalities. The model needs to perform cross-modal matching between the parsed query intent and attributes, associating them with the multimodal attribute representations of items. As shown in Fig. 1(b), the query requirement ”yellow” is absent in the textual description and necessitates extracting clues from visual information. (3) Rule-Aware Ability. Discrimination relies on quantifiable and reproducible criteria. Building upon attribute matching, the model must make judgments in accordance with established rules. As illustrated in Fig. 1(c), when a query searches for ”LV bags” while the item is a second-hand bag, this scenario cannot be resolved through attribute matching alone and requires adherence to predefined rule definitions.
@@
To model the aforementioned capabilities, we introduce LORE 3.0(Large Generative Model for Relevance), our most recent iteration of a systematic LLM practice framework for relevance tasks(the lower part of Fig 1). Firstly, We conducted necessary preliminary explorations to determine the optimal base model, features, and prompts, thereby establishing a foundation for subsequent training. Secondly, we propose a two-stage training paradigm encompassing SFT and Reinforcement Learning(RL). In the SFT stage, we progressively synthesize complete reasoning chains incorporating all processes in accordance with the two-stage relevance discrimination framework. The primary objective of this stage is to enable the model to learn reasoning patterns and elevate the upper bound of model capabilities. In the RL stage, we design Reinforcement Learning with Verifiable Rewards (RLVR)(deepseek) in conjunction with the characteristics of relevance tasks, guiding the model to align with human preferences. The objective of this stage is to prune erroneous reasoning paths while building upon the multi-dimensional reasoning capabilities the model has acquired. Thirdly, based on the identified core capability requirements, we purposefully constructed a comprehensive and challenging benchmark to systematically assess and validate the model’s performance across capability dimensions. Finally, we comprehensively transferred LLM capabilities to the online system from both the model and policy perspectives. Based on query stratification, we designed differentiated capability transfer strategies, cumulatively achieving a 27% GoodRate improvement. In summary, the main contributions of this report are as follows:
  • We deconstructed the relevance task and summarized the required model capabilities based on the deconstruction results, providing clear and comprehensive guidance for subsequent iterations of relevance models in the industry.
  • We propose a complete methodology and documented practice that encompasses the entire pipeline from preliminary work such as model selection, to training paradigm, to capability evaluation.
  • Experimentally, LORE achieved state-of-the-art results on offline datasets. Furthermore, through its subsequent empowerment of online applications, we obtained a substantial total improvement of 10% in GoodRate across online A/B experiments.
  • During LORE’s development, we explored various SFT and RL training strategies. Key findings that may benefit LLM fine-tuning in vertical domains are summarized as follows:
@@
Key Findings during SFT 1. Data scaling law shows diminishing returns. Model performance grows rapidly initially but plateaus as data increases. (Section 3.3.3). 2. More features beat fewer. Providing stable and relevant information, even if redundant, can yield gains for the model(Section 3.2.1). 3. A concise prompt with essential information works best. Information-rich but verbose prompts do not necessarily yield improvements, whereas prompts with insufficient key information lead to degradation in model capability(Section 3.2.3). 4. Naive teacher CoT distillation results in negative effects. Direct teacher CoT distillation underperforms vanilla SFT due to distribution shift. (Section 6.1)
@@
Key Findings during RL 1. Curriculum learning still works. Curriculum learning strategies based on difficulty grading significantly outperform randomly ordered strategies.(Section 3.4.3) 2. Long CoT is not necessary for better performance. No significant increase in model output length was observed during the RL process, indicating that long CoT are merely a potential byproduct when the model’s capability improves.(Section 6.2) 3. Entropy collapse rate caps model performance. In the early stages of training, the model rapidly trades entropy for performance gains. When entropy decreases to a certain level, the model’s exploration capability becomes constrained, and subsequent room for improvement diminishes.(Section 3.4.3) 4. Smart strategies effectively slow entropy collapse. The report discusses several optimization methods including clip-higher, on-policy, and entropy loss, ultimately finding that clip-higher achieves optimal performance.(Section 3.4.3)
@@ 2.1.1 Problem Formulation
From a functional perspective, e-commerce search can be regarded as the process of mapping user queries to product items. Relevance modeling refines this mapping by introducing relevance as a quantitative measure that captures the degree of match between a given query and item under both objective (e.g., attribute and constraint satisfaction) and subjective (e.g., user intent and preference) criteria. Formally, as illustrated in Fig 2, given a query $Q_{i}$ and an item $I_{i}$, the model performs pointwise discrimination to determine whether the pair is relevant or irrelevant. In its simplest form, this can be viewed as a binary classification problem.
@@ 2.2.1 Item Understanding
  • Inherent attributes denote objectively existing and reliably observable properties of the item. They are usually direct consequences of the product design or the physical carrier, such as category, shop, brand, IP, model, specification, material, color, price, year, and promotion. For such attributes, relevance tends to be close to a binary decision (relevant vs. not relevant), and the decision rules are relatively clear. Within inherent attributes, we can further distinguish a subset of characteristic attributes, which are the most central and discriminative attributes, e.g., brand, model, specification, IP, and shop name. On e-commerce platforms, these attributes are typically expressed in a clear, objective, and relatively standardized form. Consequently, for query–item matching that includes characteristic attributes, simple term-matching strategies often suffice for effective relevance judgment.
  • Derived attributes are induced from inherent attributes and usually correspond to product goals or usage effects. These attributes are not always directly observable, and relevance judgment relies more heavily on semantic understanding and reasoning, often involving subjectivity and degrees of satisfaction (e.g., “more suitable”, “more comfortable”, “more slimming”). This makes the decision problem substantially more complex. A major challenge is that different industries and categories develop their own attribute systems and surface forms, which are highly diverse and difficult to enumerate exhaustively.
@@
Presentation modality describes how attributes are exposed to the system. From this perspective, item information can be broadly divided into text-only and multi-modal (text+image) representations. The modality determines whether item understanding must rely on multi-modal modeling, and it interacts with the nature of attributes. For some attributes (e.g., color, style, material), image information is often indispensable; for others (e.g., brand, price, year), text is usually sufficient for accurate identification and judgment. In multi-modal scenarios, different attributes depend on visual information to different degrees. When characteristic attributes are expressed consistently in both text and images, multi-modal capabilities can further improve robustness. For highly subjective derived attributes (e.g., style, effect), joint understanding of text and images becomes critical for high-quality relevance judgment.
@@ 2.2.2 Query Understanding
Compared to structured item attributes, user queries express their needs in a more free-form and diverse manner. The main challenge of query understanding is to fully parse and accurately model these needs. We decompose query understanding into two core sub-tasks: entity recognition and intent recognition (entity-relation recognition). We start from single-dimensional needs and then extend to multi-dimensional needs.
(1) Entity recognition: dependence on knowledge. The goal of entity recognition is to identify all potential need points (attributes or item entities) from the query. From the perspective of difficulty and knowledge dependence, we can roughly distinguish the following scenarios:
  • Knowledge-free scenarios. Need entities can be directly extracted from surface forms, with simple syntax and semantics, and almost no background knowledge is required.
  • General world knowledge scenarios. Here, general semantic understanding is still effective, but correct recognition requires sufficient world knowledge. For example, in “pearl white” or “ivory white”, the core need is a specific shade of white, not “pearl” or “ivory” themselves; in “phone gas station”, the underlying need is a power bank, and “gas station” works as a metaphor. In such cases, higher-level semantic understanding and knowledge interact tightly, and treating the problem as pure semantic parsing is insufficient.
  • Domain-specific knowledge scenarios. These needs involve niche characteristic attributes or derived attributes that cannot be inferred from surface forms alone and must rely on structured domain knowledge. Typical examples include domain-specific jargon in vertical e-commerce, shop/brand taxonomies, model coding rules, and category-specific descriptions of functions or effects (e.g., mouthfeel descriptions for tea or wine).
  • Other challenges: misspellings and aliases. Users may express their needs using typos, homophones, aliases, or industry slang. This requires stronger integration of general and domain knowledge, and the model needs robust capabilities in spelling correction, variant normalization, and alias alignment.
@@
(2) Intent recognition (entity-relation recognition): dependence on reasoning. Intent recognition aims to determine whether the query is asking for items themselves or for item-related information (e.g., reviews, comparisons, usage guides), and to distinguish which entities are the main targets versus which are modifiers or constraints. This process depends heavily on reasoning at both the methodological and implementation levels. The main challenges include:
  • Accurately identifying the purchase purpose and item usage, thereby locating the true target of the need and distinguishing whether the query is describing derived attributes or directly referring to specific items. For example, in “charger/headphones for iPhone 17”, the actual target items are the charger or headphones, while “iPhone 17” only imposes compatibility or model constraints, corresponding to a compositional need over functionality/effect and model.
@@ 2.2.3 Path Modeling from Query to Item
In the overall framework, relevance judgment is abstracted as a process of constructing and executing paths from a query to an item. The system first maps the user query into structured retrieval and constraint paths, and then performs candidate-item filtering and judgment along these paths. From this perspective, relevance judgment decomposes into two interrelated sub-tasks: (i) constructing correct and complete paths in the semantic and attribute space, which involves systematic modeling of query intent, constraints, and their mappings to item attributes; and (ii) executing stable and reliable item judgment along these established paths to ensure consistent and reproducible assessment. The technical challenges primarily stem from the complexity and uncertainty inherent in these two processes. Below, we use the query “Summer Mint Mambo Slimming Set” to illustrate the modeling process.

### v2

@@ Abstract
Achievement. We introduce LORE(Large Generative Model for Search Relevance), a complete and sustainable framework of iterative practices for large language models(LLMs) in e-commerce search relevance, which achieves a cumulative +27% improvement on the online GoodRate metric. Over the past three years, this project has demonstrated significant improvements in relevance judgment and has undergone three full-scale iterations across key dimensions, including data, features, training paradigms, evaluation, and application. Throughout the iterative development of LORE, we have gained valuable experience and insights that we believe are worth sharing with the community in this report.
@@
Second, guided by our structural analysis, we propose a sophisticated two-stage training paradigm. In the first stage, we use progressive CoT synthesis and Supervised Fine-Tuning (SFT) to instill comprehensive capabilities. In the second, a carefully designed Reinforcement Learning (RL) phase aligns the model with human preferences. We also share key insights from our exploration of these training strategies.
@@
Finally, to overcome the challenges of real-time computation, we designed a query-based stratified deployment strategy that comprehensively transfers the offline LLM’s ability to online system, leading to substantial online performance gains.
LORE serves as both a practical solution for developing advanced e-commerce relevance systems and a methodological reference for domain-specific post-training, with insights generalizable across vertical industries.
@@ 1 Introduction
Recent advances(walmartllm; LREF; JD2; TaoSR1) have progressively transitioned the post-training paradigm from rudimentary classification-oriented Supervised Fine-Tuning (SFT) to sophisticated enhancement of reasoning capabilities. While existing work has modeled the Chain-of-Thought (CoT)(CoT) processes of LLMs from various perspectives to enhance their capabilities, significant blind spots remain. For example, ELLM(ELLM) conceptualizes relevance as a process of attribute extraction and matching between queries and items. However, this approach falters in boundary cases that demand rule-based judgment. For instance, Lacking fine-grained rules for color discrimination, it might incorrectly match a ”lake blue top” with a ”sky blue top”. To address this, subsequent approach like LREF(LREF) and TaoSR1(TaoSR1) advance beyond simple attribute matching by explicitly modeling rule-following CoT, thereby achieving rule-awareness. Nevertheless, their reliance solely on textual information introduces a significant limitation: the inability to process visual cues. As a result, given the query ”blue top,” the models cannot confirm a match if an item’s title omits the color, even its image clearly displays a blue top. These limitations ultimately stem from a failure to address a fundamental question that defines the capabilities required for relevance assessment: What constitutes a comprehensive CoT representation for such tasks?
@@
To systematically address this question, we first deconstruct the relevance assessment, as detailed in Section 2. Our analysis breaks this process down into two fundamental steps: (1) Path Construction: mapping query requirements to specific attributes and establishing semantic paths between queries and products; (2) Path Following: executing precise judgments along this path on an attribute-by-attribute basis. This two-stage decomposition allows us to pinpoint the essential capabilities that model must possess, which we summarize as follows: (1) Knowledge and Reasoning(deepseek; skywork; kimi1.5; knowledge1; knowledge2). Knowledge is required to decipher domain-specific terminology in queries (e.g., ”morning C, evening A” as Vitamin C and retinol) and interpret ambiguous item details. Reasoning then leverages this knowledge to infer higher-level user intent (e.g., deducing that ”for mom” implies a middle-aged demographic), as illustrated in Fig. 1(a). (2) Multi-modal Matching(multi-modal1; multi-modal2; multi-modal3). Since item attributes are often distributed across various modalities, this ability is crucial. The model must perform cross-modal alignment, connecting textual query requirements with visual evidence from the item. As shown in Fig. 1(b), a query for ”yellow” must be validated against the item’s image when the color is not mentioned in its textual description. (3) Rule-Aware. Rules serve as the concrete embodiment of human subjective judgment, supplementing general logic by providing explicit, quantifiable, and reproducible criteria for relevance assessment. it must adhere to these predefined rules to make nuanced judgments. For instance, as depicted in Fig. 1(c), classifying a ”second-hand LV bag” against a query for ”LV bags” is not merely an attribute-matching problem. It requires applying a specific rule about item condition (e.g., new vs. used) to ensure accuracy.
@@
To instill these capabilities, we introduce LORE 3.0(Large Generative Model for Search Relevance), our most recent iteration of the systematic LLM practice framework for relevance task, as depicted in the lower part of Fig 1. Firstly, we establish a robust foundation by conducting preliminary explorations to identify the optimal base model, feature set, and prompt structure. Secondly, building on this foundation, we propose a two-stage training paradigm encompassing SFT and Reinforcement Learning(RL). The SFT stage is designed to instill comprehensive reasoning abilities. Following our two-step discrimination framework, we synthesize reasoning trajectories through a progressive synthesis strategy and elevate the model’s capability upper bound via distillation. The subsequent RL stage aims to align the model with human preferences. To this end, we adapt Reinforcement Learning with Verifiable Rewards (RLVR)(deepseek) and tailor it to the nuances of relevance tasks. This process prunes flawed reasoning paths while preserving the rich, multi-dimensional capabilities established during SFT. Thirdly, to rigorously evaluate our model, we constructed a comprehensive and challenging benchmark designed specifically around the core capabilities identified earlier. This allows for a systematic assessment of the model’s performance across each dimension. Finally, we implemented a practical strategy to transfer these enhanced LLM capabilities into our online system. By stratifying queries and applying differentiated deployment policies, we achieved a cumulative 27% improvement in GoodRate. In summary, the main contributions of this work are as follows:
  • We present a theoretical deconstruction of the relevance task, identifying a core set of model capabilities. This analysis establishes a principled blueprint for designing future relevance models.
  • Building on the theory, we introduce LORE, a replicable, end-to-end methodology for the entire model lifecycle—from data synthesis and training to online deployment. It serves as a comprehensive guide for industrial practice.
  • We demonstrate LORE’s effectiveness through comprehensive offline and online validation. The framework not only achieves SOTA performance on offline benchmarks but also delivers a substantial 27% cumulative improvement in online GoodRate upon deployment.
  • We distill key insights and best practices from our extensive exploration of SFT and RL strategies. These findings provide actionable guidance for fine-tuning LLMs in other specialized domains:
@@
Key Findings during SFT 1. More features beat fewer. Providing stable and relevant information, even if redundant, can yield gains for the model(Section 3.2.1). 2. A concise prompt with essential information works best. Information-rich but verbose prompts do not necessarily yield improvements, whereas prompts with insufficient key information lead to degradation in model capability(Section 3.2.3). 3. Data scaling law shows diminishing returns. Model performance grows rapidly initially but plateaus as data increases. (Section 3.3.3). 4. Naive teacher CoT distillation results in negative effects. Direct teacher CoT distillation underperforms vanilla SFT due to distribution shift. (Section 6.1)
@@
Key Findings during RL 1. Curriculum learning still works. Curriculum learning strategies based on difficulty grading significantly outperform randomly ordered strategies.(Section 3.4.3) 2. Entropy collapse rate caps model performance. In the early stages of training, the model rapidly trades entropy for performance gains. When entropy decreases to a certain level, the model’s exploration capability becomes constrained, and subsequent room for improvement diminishes.(Section 3.4.3) 3. Smart strategies effectively slow entropy collapse. The report discusses several optimization methods including clip-higher, on-policy, and entropy loss, ultimately finding that clip-higher achieves optimal performance.(Section 3.4.3) 2. Long CoT is not necessary for better performance. No significant increase in model output length was observed during the RL process, indicating that long CoT are merely a potential byproduct when the model’s capability improves.(Section 6.2)
@@ 2.1.1 Problem Formulation
At its core, e-commerce search can be conceptualized as a mapping from user queries to product items. The role of relevance modeling is to refine this mapping by establishing a quantitative measure—relevance—that evaluates the quality of this match. This measure accounts for both objective criteria (e.g., attribute and constraint satisfaction) and subjective factors (e.g., user intent and preference). Formally, as illustrated in Fig 2, given a query $Q_{i}$ and an item $I_{i}$, the model performs pointwise discrimination to determine whether the pair is relevant or irrelevant. In its simplest form, this can be viewed as a binary classification problem.
@@ 2.2.1 Item Understanding
  • Inherent attributes denote objectively existing and reliably observable properties of the item. They are usually direct consequences of the product design or the physical carrier, such as category, shop, brand, IP, model, specification, material, color, price, year, and promotion. For such attributes, relevance tends to be close to a binary decision (relevant vs. not relevant), and the decision rules are relatively clear. Within inherent attributes, we can further distinguish a subset of characteristic attributes, which are the most central and discriminative attributes, e.g., brand, model, IP, and shop name. On e-commerce platforms, these attributes are typically expressed in a clear, objective, and relatively standardized form. Consequently, for query–item matching that includes characteristic attributes, simple term-matching strategies often suffice for effective relevance judgment.
  • Derived attributes are induced from inherent attributes and usually correspond to product goals or usage effects. These attributes are not always directly observable, and relevance judgment relies more heavily on semantic understanding and reasoning, often involving subjectivity and degrees of satisfaction (e.g., “style”, “ audience”, “function”). This makes the decision problem substantially more complex. A major challenge is that different industries and categories develop their own attribute systems and surface forms, which are highly diverse and difficult to enumerate exhaustively.
@@
Presentation modality describes how attributes are exposed to the user. From this perspective, item information can be broadly divided into text-only and multi-modal (text+image) representations. The modality determines whether item understanding must rely on multi-modal modeling, and it interacts with the nature of attributes. For some attributes (e.g., color, style, material), image information is often indispensable; for others (e.g., brand, price, year), text is usually sufficient for accurate identification and judgment. In multi-modal scenarios, different attributes depend on visual information to different degrees. When characteristic attributes are expressed consistently in both text and images, multi-modal capabilities can further improve robustness. For highly subjective derived attributes (e.g., style, effect), joint understanding of text and images becomes critical for high-quality relevance judgment.
@@ 2.2.2 Query Understanding
In contrast to the structured attributes of items, user queries are often free-form and express a diverse range of needs. The primary challenge in Query Understanding, therefore, is to comprehensively parse and accurately model these expressions. To address this, we decompose Query Understanding into two core sub-tasks: Entity Recognition and Intent Recognition (entity-relation recognition). Our analysis begins with single-dimensional queries and progressively extends to more complex, multi-dimensional ones.
(1) Entity recognition. The objective of Entity Recognition is to extract all entities from a query that represent user needs, such as specific product attributes or item types. Based on their inherent difficulty and dependence on domain knowledge, these recognition tasks can be broadly categorized into the following scenarios:
  • Knowledge-free scenarios. the entities represented users need can be directly extracted from surface forms, with simple syntax and semantics, and almost no background knowledge is required.
  • General world knowledge scenarios. While general semantic understanding is foundational, accurate interpretation in these cases hinges on substantial world knowledge. For instance, in “pearl white” or “ivory whit”, the user’s intent is a specific shade of white, not the literal objects “pearl” or “ivory”. Similarly, “phone gas station” is a metaphorical expression for a “power bank”, where “gas station” signifies a source of energy. Such examples demonstrate that higher-level semantic reasoning and world knowledge are tightly intertwined, rendering an approach based solely on pure semantic parsing insufficient.
  • Domain-specific knowledge scenarios. These requirements involve niche characteristic attributes or derived attributes that cannot be inferred from surface forms alone and must rely on structured domain knowledge. Typical examples include domain-specific jargon in vertical e-commerce, shop/brand taxonomies, model coding rules, and category-specific descriptions of functions or effects (e.g., mouthfeel descriptions for tea or wine).
  • Other challenges: misspellings and aliases. Users often express their needs using non-standard language, such as typos, homophones, aliases, or industry slang. Effectively addressing this requires a tighter integration of general and domain knowledge, demanding that the model possess robust capabilities for spelling correction, variant normalization, and alias alignment.
@@
(2) Intent recognition Intent recognition is the task of discerning whether a query is transactional (seeking an item) or informational (seeking related content such as reviews, comparisons, or usage guides). It also involves distinguishing the core subject of the query from its modifiers and constraints. This entire process is highly dependent on sophisticated reasoning. Key challenges in this area include:
  • Accurately identifying the purchase purpose and item usage, thereby locating the true target of the need and distinguishing whether the query is describing derived attributes or directly referring to specific items. For example, in “summer mint mambo slimming set”, the target item is “set”, while the remaining modifiers serve as attribute constraints covering scenario (summer), function (slimming), brand (mambo), and style (mint).
@@ 2.2.3 Path Modeling from Query to Item
In the overall framework, relevance judgment is abstracted as a process of constructing and executing paths from a query to an item. The system first maps the user query into structured retrieval and constraint paths, then performs candidate-item filtering and judgment along these paths. From this perspective, relevance judgment decomposes into two interrelated sub-tasks: (1) constructing correct and complete paths in the semantic and attribute space, which involves systematic modeling of query intent, constraints, and their mappings to item attributes; and (2) executing stable and reliable item judgment along these established paths to ensure consistent and reproducible assessment. The technical challenges primarily stem from the complexity and uncertainty inherent in these two processes. Below, we use the query “Summer Mint Mambo Slimming Set” to illustrate the modeling process.